through:(1,2), slope=7
Answer by alicealc(293)

line equation with slope m and passes through (x1,y1):
y - y1 = m*(x - x1)

so, if the line passes through (1, 2) and has the slope of 7:
y - 2 = 7*(x - 1)
y - 2= 7x - 7
y = 7x - 7 + 2
y = 7x - 5